OpenStudy (nuccioreggie):

Nancy left a bin outside in her garden to collect rainwater. She notices that 1 over 2 gallon of water fills 2 over 3 of the bin. Write and solve an equation to find the amount of water that will fill the entire bin. Show your work. Explain your answer in words.

OpenStudy (nuccioreggie):

@Michele_Laino Please explain

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

we have to compute this: \[\frac{1}{2} \times \frac{3}{2} = ...?\]

OpenStudy (nuccioreggie):

I think this is the answer but im not sure(1/2) / (2/3) = x / (3/3) --- 1/2 gal to 2/3 bin = x gal to 3/3 bin) cross multiply (2/3)(x) = (1/2)(3/3) 2/3x = 3/3 x = (3/3) / (2/3) x = 3/3 * 3/2 x = 9/6 = 1 1/2 the amount of water that would fill the entire bin is : 1 1/2 gallons or 1.5 gallons

OpenStudy (michele_laino):

in order to fill completely a bin, we need 3/2 times 1/2 gallon

OpenStudy (z4k4r1y4):

if 1/2 a gallon fills 2/3 of the bin. 1/4 will fill 1/3. therefor 3/4 gallons will fill the whole bin. Is that right?

OpenStudy (nuccioreggie):

Ok can u right the answer out for me explaining plea ese

OpenStudy (z4k4r1y4):

if we know how much is needed to fill 2/3. all we need to work out is how much is needed for the last 1/3. do you understand?

OpenStudy (z4k4r1y4):

now, half of 2/3 is 1/3. it takes 1/2 a gallon to fill 2/3 so half of that (1/4) will fill 1/3.

OpenStudy (nuccioreggie):

So thas the answer??

OpenStudy (z4k4r1y4):

so if 1/2 a gallon fills 2/3 and 1/4 gallon fills 1/3 of the bin, the full bin will need 1/2 + 1/4 gallons of water which = 3/4 gallons.

OpenStudy (nuccioreggie):

Thank you is he correct @Michele_Laino

OpenStudy (michele_laino):

thank you! @nuccioreggie @Z4K4R1Y4

OpenStudy (z4k4r1y4):

you're welcome.